Grade 11Discuss with Askiitians Tutors

if a,b,c are positive ,then show that (a+b)(b +c)(c+a)>=8abc

Ayushi Agrawal
13 Years agoGrade 11

1 Answer

Aman  Bansal
13 Years ago

Dear Ayushi,

by the principle of arithmeticmean>=geometricmean

(a+b)/2 >= (ab)^(0.5)------1
(c+b)/2 >= (cb)^(0.5)--------2
(a+c)/2 >= (ac)^(0.5)-----------3

as all abc are positive so by multiplying all the three we get

(a+b)(b+c)(c+a)/8 >= (abbcca)^(0.5)
(a+b)(b+c)(c+a)/8 >= abc
(a+b)(b+c)(c+a) >= 8abc

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert